Passive Seismic Imaging

Bradley W. Artman, Ph.D. Stanford University, 2007

Adviser: Professor Biondo L. Biondi

Passive seismic imaging, the process of synthesizing active surface data by correlating ambient noise records, has languished in the realm of theory since the original concept was presented 1968. Very few experimental tests have been fielded since then, and none provided convincing results. The recent surge in interest in the idea has generated a substantial literature that has proved the validity of the theory beyond the initial conjecture, as well as experimentally at the lab scale and the global scale. This thesis makes a substantive contribution to the field by extending the processing algorithms available for passive seismic data and by analyzing a real 3d data volume from a 2300 station areal array.

Direct migration of transmission wavefields, introduced herein, is a robust, convenient processing method. Any conventional shot profile migration algorithm that maintains separate up-coming and down-going wavefields can be easily modified to use passive data. The only change required is to use the data for the initial condition of the source wavefield. Processing passive data by direct migration can provide substantial computational savings. This is especially true if multiple images are produced from the total volume of data collected. However, synthesized shot gathers may be required for velocity analysis or addressing common data problems such as noise removal or aliasing.

Over 35 hours of passive data collected at the Valhall development in the North Sea was processed to produce synthesized shot gathers and migrated images. The synthesized shot gathers did not provide any subsurface information. Instead, they were dominated by a direct arrival through the water column from the gas flare at a production facility 40 km away from the array. The energy from the gas flare is a coherent noise source that arrives at the array at such a shallow angle that it is confined to the water column and does not refract into the subsurface. As a coherent noise source, the energy manufactures events in image results that are not subsurface reflectors. Removing the far-field noise from the synthesized gathers also removes the interpretable, though false, events from images.

Approved for publication:

By:

For the Geophysics Department